OpenStudy (anonymous):

How do I evaluate arccsc(sec2)?

OpenStudy (anonymous):

isn't it sec(2)?

OpenStudy (dumbcow):

oh did i misunderstand the problem? its not sec^2

OpenStudy (anonymous):

cause the question is arcsc(sec(2))

OpenStudy (anonymous):

\[\csc ^-1(\sec (2))\]

OpenStudy (dumbcow):

oh in this case use a calculator http://www.wolframalpha.com/input/?i=arccsc%28sec%282%29%29

OpenStudy (anonymous):

how do you do it by hand?

OpenStudy (anonymous):

the solution is suppose to be pi/2 - 2, but i dont know how to get it.

OpenStudy (dumbcow):

\[\theta = \csc ^{-1} (\sec 2)\] \[\csc \theta = \sec 2\] take reciprocal \[\sin \theta = \cos 2\] draw triangle  from triangle \[\tan (\frac{\pi}{2} - 2) = \frac{\sin \theta}{\cos \theta} = \tan \theta\] therefore \[\theta = \frac{\pi}{2} - 2\]

OpenStudy (dumbcow):

did that make sense?

OpenStudy (anonymous):

i'm kind of confused on how you drawn the triangle.

OpenStudy (dumbcow):

ahh ok cos(2) = adj/hyp = sin(theta)/1 this is why hypotenuse is 1 and adjacent side is sin(theta) remember in right triangle the 2 angles must add up to 90 or pi/2 so if one angle is 2 the other must be (pi/2) -2

OpenStudy (anonymous):

Ok, I think I got it now. Thanks!

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

I might need your help again later :)